# SceneCOT: Eliciting Grounded Chain-of-Thought Reasoning in 3D Scenes

### v2

@@ 5.4 Reasoning Chain Visualization
Finally, we present a case study in Fig.˜5 to illustrate the strengths and limitations of our framework. In the first example, the model grounds the doors at the agent’s 2 o’clock and constructs the corresponding visual clue obj_prob. Although the predicted labels differ slightly from the ground-truth “door,” the accurate probabilities and semantic similarity allow the model to generate the correct answer. In the second example, the model identifies all monitors and uses the symbolic engine to compute their relative positions with 2D polar coordinates. By integrating spatial information with object probabilities, the model infers the answer “4 o’clock.” These two cases demonstrate interpretable reasoning chains for challenging tasks, making error sources easier to diagnose. In the third example, however, the model fails in the final reasoning step despite an accurate visual clue, echoing the upper-bound analysis in Tab.˜3 where Navigation performance peaks at 87.2. This highlights a gap due to limited reasoning capability in the base model, which could be improved as foundation models like multimodal LLM advance. Overall, these visualizations showcase both the progress and remaining challenges of SceneCOT, pointing toward more robust and interpretable 3D reasoning frameworks.
@@ 6 Conclusion
We presented SceneCOT, a framework for step-by-step grounded reasoning in 3D scenes. Unlike prior approaches that treat 3D question answering as a single-step task, SceneCOT decomposes reasoning into task recognition, region localization, entity grounding, and grounded reasoning, closely reflecting the human problem-solving process. To enable this, we built SceneCOT-185K, the first large-scale dataset of grounded Chain-of-Thought reasoning traces with rich stepwise annotations linking language and 3D context. Experiments show that SceneCOT achieves strong performance on both situated reasoning benchmarks and the Beacon3D benchmark, where it significantly improves grounding–QA coherence compared to object-centric baselines. Beyond accuracy, our method generates interpretable reasoning traces that make the decision process transparent. These results demonstrate that step-by-step grounded reasoning is both effective and necessary for robust 3D scene understanding, laying a foundation for advancing multimodal LLMs toward human-like reasoning in real-world 3D environments.
Appendix A SceneCOT
@@ A.2.2 Visual Clue Construction
In the Grounded Reasoning step, the model generates the final answer by integrating both semantic information and a visual clue derived from object probabilities. We outline the algorithm for constructing this visual clue during the reasoning process. Given a prefix sequence (i.e., the input_ids of the text prompt), the algorithm returns a modified sequence that incorporates relevant visual clues based on object probabilities, spatial locations, and semantic labels. This enhanced sequence is then fed back into the LLM to generate the final response. The component functions—build_obj_prob, build_obj_loc_prob, and build_obj_loc_plr_prob—are illustrated in Fig.˜15 and Fig.˜3. Additionally, Fig.˜13 provides a visual comparison between the 2D polar coordinate system and the 3D spatial coordinate system used in our framework.
@@ A.3 Implementation Details of the 3D Visual Grounding Module
We illustrate the implementation of 3D Visual Grounding Module in Fig.˜14. In the framework, the object feature is extracted by PQ3D, while the text embedding is extracted by the embedding tokenizer based on the grounding text. We train PQ3D based on the grounding data in Jia et al. (2024), including the scenes in ScanNet, 3RScan, and MultiScan.
@@ B.1 Data Generation Pipeline of SceneCOT-185K
Our data generation pipeline is shared between Object-Centric Reasoning and Situated Reasoning, with Object-Centric Reasoning formulated as an Attribute sub-task under the broader Situated Reasoning framework. The overall process is illustrated in Fig.˜15.
@@ B.2 Data Generation Details and Examples of GQA3D
To enable the training of Object-Centric Reasoning in non-situated scenarios, we incorporate GQA3D as a key component of the overall dataset. The primary distinction from Situated Reasoning lies in the nature of the grounding text, which describes the target object from a global perspective rather than an egocentric view. Additionally, this task does not involve multi-object grounding or reasoning. We use GPT-4o to generate diverse question-answer pairs based on a given object image and its corresponding grounding text. We provide the system prompt in Fig.˜16.
@@ B.3 Data Quality
We conducted a manual quality check to verify the data quality. First, we manually checked the object IDs in MSQA. Note that all object IDs outside the scene were filtered out in our dataset. Since MSQA is based on situated scene graphs, all objects in QA pairs correspond to specific sub-regions, as ensured by the MSQA paper. We randomly sampled 50 examples each from counting, spatial relationship, refer, attribute, and navigation tasks, totaling 250 examples. Annotators were given the question, answer, and object IDs to verify if the IDs align with the QA pairs. For example, if the question is “How many chairs are on my left?” with the answer “three,” but the target object IDs are ["chair-1", "chair-3"], then the annotation is incorrect. The results are shown in Tab.˜7. The results demonstrate the reliability of our annotation process.
Next, we verify the quality of the generated QA pairs in GQA3D. GQA3D uses Nr3D as its data source to create QA pairs with reasoning traces. The grounding text is mostly human-annotated, ensuring data quality. We generated new QA pairs based on object images using GPT-4V. To ensure quality, we randomly sampled 100 QA pairs and manually verified their correctness, providing both the object images and QA pairs for review. The results are shown in Tab.˜8. The results confirm the reliability of our data. The strong performance of Chat-Scene and SceneCOT further highlights the value of our dataset.
@@ C.1 Additional Baseline Comparison
| SceneCOT | ours | ✓ | 47.9 | 82.1 | 49.6 | 47.2 | 51.6 | 80.3 | 55.6 | 58.9 | 23.2 |
In our main results, we focus on object-centric methods such as Chat-Scene, LEO, and MSR3D to clearly analyze both QA performance and grounding–QA coherence. To broaden the comparison, we also include LLaVA-3D as a strong voxel-based baseline. As shown in Tab.˜9, SceneCOT achieves competitive performance on both Situated Reasoning and Object-Centric Reasoning. An interesting direction for future work is integrating SceneCOT with voxel-based approaches.
C.2 Additional Reasoning Chain Visualization Results
@@
Appendix D Limitations
@@
First, our framework focuses on the tasks pre-defined in MSQA, which is limited in more complex scenarios such as embodied AI. For example, we do not consider the long-horizon tasks such as embodied task planning. Recently, SG3D Zhang et al. (2024c) proposes a new benchmark to evaluate 3D-VL models’ capabilities of grounded task planning in embodied scenarios. We will consider extending our framework to this task in the future.
@@
Appendix E Broader Impact
@@
Appendix F LLM Usage Statement

### v3

@@ 5.5 Reasoning Chain Visualization
Finally, we present a case study in Fig.˜6 to illustrate the strengths and limitations of our framework. In the first example, the model grounds the doors at the agent’s 2 o’clock and constructs the corresponding visual clue obj_prob. Although the predicted labels differ slightly from the ground-truth “door,” the accurate probabilities and semantic similarity allow the model to generate the correct answer. In the second example, the model identifies all monitors and uses the symbolic engine to compute their relative positions with 2D polar coordinates. By integrating spatial information with object probabilities, the model infers the answer “4 o’clock.” These two cases demonstrate interpretable reasoning chains for challenging tasks, making error sources easier to diagnose. In the third example, however, the model fails in the final reasoning step despite an accurate visual clue, echoing the upper-bound analysis in Tab.˜3 where Navigation performance peaks at 87.2. This highlights a gap due to limited reasoning capability in the base model, which could be improved as foundation models like multimodal LLM advance. Overall, these visualizations showcase both the progress and remaining challenges of SceneCOT, pointing toward more robust and interpretable 3D reasoning frameworks.
@@ 6 Conclusion
We presented SceneCOT, a framework for step-by-step grounded reasoning in 3D scenes. Unlike prior approaches that treat 3D question answering as a single-step task, SceneCOT decomposes reasoning into task recognition, region localization, entity grounding, and grounded reasoning, closely reflecting the human problem-solving process. To enable this, we built SceneCOT-185K, the first large-scale dataset of grounded Chain-of-Thought reasoning traces with rich stepwise annotations linking language and 3D context. Experiments show that SceneCOT achieves strong performance on both situated reasoning benchmarks and the Beacon3D benchmark, where it significantly improves grounding-QA coherence compared to object-centric baselines. Beyond accuracy, our method generates interpretable reasoning traces that make the decision process transparent. These results demonstrate that step-by-step grounded reasoning is both effective and necessary for robust 3D scene understanding, laying a foundation for advancing multimodal LLMs toward human-like reasoning in real-world 3D environments.
Reproducibility statement
In this paper, we propose a new reasoning framework together with a large-scale dataset. We present the implementation details of the model design and training in both the main paper and the appendix. In Sec.˜3.1, we define the different types of thoughts based on question categories. In Sec.˜3.2, we describe how the modular components are trained and the pre-trained models we adopt. Hyper-parameters used for training and evaluation are provided in Sec.˜A.5. Additional details include the definition of the symbolic engine in Sec.˜A.2, the visual clue construction algorithm in Fig.˜15, and the design of the 3D grounding module in Sec.˜A.3.
Ethics Statement
This work adheres to the ICLR Code of Ethics.††https://iclr.cc/public/CodeOfEthics Our study focuses on improving 3D scene reasoning through step-by-step grounded reasoning. We use only publicly available 3D vision-language datasets, none of which contain personal or sensitive information. No human subjects were involved in this research.
Our method inherits biases and limitations from pre-trained models and datasets, such as potential semantic or linguistic biases in object labels and answers. While our framework improves interpretability and grounding-QA coherence, it is not intended for direct deployment in safety-critical domains (e.g., autonomous navigation or surveillance) without further safeguards. Code and processed data will be released to support transparency and reproducibility.
Appendix A SceneCOT Framework
@@ A.2.2 Visual Clue Construction
In the Grounded Reasoning step, the model generates the final answer by integrating both semantic information and a visual clue derived from object probabilities. We outline the algorithm for constructing this visual clue during the reasoning process. Given a prefix sequence (i.e., the input_ids of the text prompt), the algorithm returns a modified sequence that incorporates relevant visual clues based on object probabilities, spatial locations, and semantic labels. This enhanced sequence is then fed back into the LLM to generate the final response. The component functions—build_obj_prob, build_obj_loc_prob, and build_obj_loc_plr_prob—are illustrated in Fig.˜16 and Fig.˜3. Additionally, Fig.˜14 provides a visual comparison between the 2D polar coordinate system and the 3D spatial coordinate system used in our framework.
@@ A.3 Implementation Details of the 3D Visual Grounding Module
We illustrate the implementation of 3D Visual Grounding Module in Fig.˜15. In the framework, the object feature is extracted by PQ3D, while the text embedding is extracted by the embedding tokenizer based on the grounding text. We train PQ3D based on the grounding data in Jia et al. (2024), including the scenes in ScanNet, 3RScan, and MultiScan.
@@ B.1 Data Generation Pipeline of SceneCOT-185K
Our data generation pipeline is shared between Object-Centric Reasoning and Situated Reasoning, with Object-Centric Reasoning formulated as an Attribute sub-task under the broader Situated Reasoning framework. The overall process is illustrated in Fig.˜16.
@@ B.2 Data Generation Details and Examples of GQA3D
To enable the training of Object-Centric Reasoning in non-situated scenarios, we incorporate GQA3D as a key component of the overall dataset. The primary distinction from Situated Reasoning lies in the nature of the grounding text, which describes the target object from a global perspective rather than an egocentric view. Additionally, this task does not involve multi-object grounding or reasoning. We use GPT-4o to generate diverse question-answer pairs based on a given object image and its corresponding grounding text. We provide the system prompt in Fig.˜17.
@@ B.3 Data Quality
We conducted a manual quality check to verify the data quality. First, we manually checked the object IDs in MSQA. Note that all object IDs outside the scene were filtered out in our dataset. Since MSQA is based on situated scene graphs, all objects in QA pairs correspond to specific sub-regions, as ensured by the MSQA paper. We randomly sampled 50 examples each from counting, spatial relationship, refer, attribute, and navigation tasks, totaling 250 examples. Annotators were given the question, answer, and object IDs to verify if the IDs align with the QA pairs. For example, if the question is “How many chairs are on my left?” with the answer “three,” but the target object IDs are ["chair-1", "chair-3"], then the annotation is incorrect. The results are shown in Tab.˜8. The results demonstrate the reliability of our annotation process.
Next, we verify the quality of the generated QA pairs in GQA3D. GQA3D uses Nr3D as its data source to create QA pairs with reasoning traces. The grounding text is mostly human-annotated, ensuring data quality. We generated new QA pairs based on object images using GPT-4V. To ensure quality, we randomly sampled 100 QA pairs and manually verified their correctness, providing both the object images and QA pairs for review. The results are shown in Tab.˜9. The results confirm the reliability of our data. The strong performance of Chat-Scene and SceneCOT further highlights the value of our dataset.
@@ C.1 Additional Baseline Comparison
| SceneCOT | ours | ✓ | 47.7 | 82.1 | 49.6 | 47.2 | 51.6 | 80.3 | 55.6 | 58.9 | 23.2 |
In our main results, we focus on object-centric methods such as Chat-Scene, LEO, and MSR3D to clearly analyze both QA performance and grounding-QA coherence. To broaden the comparison, we also include LLaVA-3D as a strong voxel-based baseline. As shown in Tab.˜10, SceneCOT achieves competitive performance on both Situated Reasoning and Object-Centric Reasoning. An interesting direction for future work is integrating SceneCOT with voxel-based approaches.
C.2 Additional Evaluation and Baseline Comparison
| Method | SQA3D (EM-R) | SQA3D-G (F1@50) | ScanQA (EM-R) | ScanQA-G (F1@50) |
| Chat-Scene | 36.0 | 3.4* | 24.5 | 4.1* |
| LEO | 36.7 | – | 20.04 | – |
| SceneCOT | 39.7 | 51.6 | 21.0 | 40.8 |
Note on Chat-Scene: Despite following official instructions and using 8-GPU reproduction attempts, grounding performance remained low. Reports (e.g., GitHub Issue #38) suggest the model is highly sensitive to training settings.
We assess our model on SQA3D and ScanQA under zero-shot settings and compare it with Chat-Scene and LEO. Both Chat-Scene and LEO are fine-tuned on SceneCOT-QA; specifically, LEO is fine-tuned on the SceneCOT-QA dataset, while Chat-Scene is fine-tuned on the SceneCOT-grounded QA variant, where the model must first predict the target object IDs before generating the answer. The results are summarized in Table 11.
We evaluate both QA and grounding performance. The results indicate that our QA performance is comparable to Chat-Scene, while our grounding performance is substantially stronger. This reinforces our claim that although SceneCOT does not rely on scene tokens, it explicitly grounds objects before answering. Since answers must be inferred solely from grounded results, the model demonstrates strong grounding–QA coherence, whereas Chat-Scene shows weak coherence between its QA performance and grounding performance.
C.3 Details of QA-driven Grounding Benchmarks
| # Instance | Accuracy |
| 33 | 93.3% |
In Sec.˜C.2, we evaluate the grounding performance of the models on ScanQA and SQA3D. For ScanQA, we utilize the object instance IDs provided by the official annotations. For SQA3D, we built an agent workflow and chose the rules defined in Sec.˜A.2 to extract the question-answer related object IDs. We denote the corresponding data as “ScanQA-G” and “SQA3D-G”. To guarantee the data quality, we conducted manual quality verification. The results is shown in Tab.˜12. The result reveals the reliability of SQA3D-G.
C.4 Additional Reasoning Chain Visualization Results
@@
Appendix D Further Analysis and Efficiency
D.1 Inference Latency and Trade-offs
Our framework primarily focuses on enabling a reliable and transparent reasoning process, which requires incorporating all necessary contextual information during inference. This approach inevitably incurs higher inference time due to multiple module calls and the construction of detailed reasoning contexts. We provide a qualitative comparison of latency and token length in Table 13.
| Method | Object instance information? | Reasoning token length | Inference latency |
| LEO | No | 0 | 4.8s |
| LLaVA-3D | No | 0 | 0.2s |
| Chat-Scene | No | 0 | 0.5s |
| SceneCOT | Yes | 350–1500 | 10.4s |
The majority of the latency in SceneCOT (10.4s total) stems from LLM sequence generation, specifically Stage 1–3 generation (4.2s) and Stage 4 generation (3.8s), while visual clue construction remains relatively moderate at 2.4s. We believe this represents a reasonable trade-off between transparent, interpretable reasoning and inference speed.
D.2 Accuracy of Task and Region Recognition
In our framework, errors from Task Recognition and Region Localization can be largely ignored due to their high accuracy. As shown in Table 14, both sub-tasks achieve extremely high performance, which we attribute to the strong commonsense priors inherent in large language models.
| Module | Accuracy (%) |
| Question type recognition | 99.4 |
| Region recognition | 100.0 |
For Region Localization, we employ a rule-based filtering mechanism. For instance, given an instruction like “Now I need to ground the objects on my right,” the system applies predefined rules to filter objects within the corresponding sub-region, ensuring accurate object selection aligned with the model’s recognition results.
Appendix E Limitations
@@
First, our framework focuses on the tasks pre-defined in MSQA, which is limited in more complex scenarios such as embodied AI. For example, we do not consider the long-horizon tasks such as embodied task planning. Recently, SG3D Zhang et al. (2024b) proposes a new benchmark to evaluate 3D-VL models’ capabilities of grounded task planning in embodied scenarios. We will consider extending our framework to this task in the future.
@@
Appendix F Broader Impact
@@
Appendix G LLM Usage Statement